# 1]FAIR at Meta

### v1

@@ 5.3 Timings
The FLOPs necessary to compute the sparse attention are about 20$\times$ lower than for the full attention. The reduction of runtime is 65%, because the computation is not as well optimized as FlashAttention and includes overheads.
5.4 Comparison with the state of the art
For comparison with the SOTA method RetrievalAttention Liu et al. (2024a), we run Saap on a Llama 3 model fine-tuned by Gradient AI to support 262k contexts. We set $\ell=32$, yielding a selectivity around 5% for textual content.
Table 3 shows that the baseline numbers are often different, probably because of slightly different experimental settings (prompts). However, the gap between Saap and the full attention baseline is very similar to the RetrievalAttention results. Note that in our experimental setting, the dense part of the attention is 1+2047 tokens, while it is 128+512 for RetrievalAttention. However, for long-range tasks, this does not matter much as is still a tiny fraction of the attention size (see the ablations in Section 5.5).
The critical advantage of Saap is that the slowest stage (training) is performed offline, while RetrievalAttention’s RoarGraph index Chen et al. (2024a) has to be built once the context is available, which takes several minutes for a 100k context. As mentioned in section 2, this makes graph-based indexing of a KV cache attractive only in settings where the same KV cache is reused multiple times. For Saap, the pre-fill of our index takes a few seconds on modern GPUs.
5.5 Ablation experiments
In this section, we run Saap with some variations on the default design, and measure their impact.
Q-model architecture and training.
In Appendix 8, we analyze some variants of the Q-model training. This analysis shows that using long-range queries-key pairs for training is useful. Using a residual architecture rather than a simple MLP does not help.
Dense context size.
Saap’s default dense context is 2048 tokens (the attention sink of size 1 plus the 2047 most recent keys), see Section 3.5. We experiment with some other settings for attention sink+dense context: 1+511, 1+63 and 128+512. The results in Appendix 8 show that the performance remains similar.
Batched queries.
Saap combines all four queries belonging to the same group into a single search, see Section 4.4. These queries could also be performed independently. In Appendix 8, we show that batching the queries corresponding to the same KV head outperforms ranking the buckets independently for each query.
6 Conclusion
We introduced Saap, a new method for non-exhaustive attention computations in LLMs. Saap is data adaptive, while incurring negligible run-time or memory overhead compared to exhaustive attention. We show the efficiency of Saap for typical long-context tasks, with significant computation speedups and a small impact on accuracy.
There are numerous more complex variants we tried which did not yield consistent improvements: increasing the model capacity, training a key classification model, fine-tuning the query classification model to the current prompt, using causal masking in the training batches, training on book summarization data rather than books to force long-range queries. In general, we observed that good results from experiments at the level of one head do not guarantee an improved e2e performance. This is problematic, because the training itself is local to one head in the current setting.
There are natural extensions to Saap that we leave as future work. Saap can be combined with vector compression to reduce the amount of GPU memory used. Currently, we apply Saap at the generation stage because this is when memory I/O has the strongest impact on performance, but it should be possible to improve pre-fill time as well.

### v2

@@ 6.3 Timings
runtime 50 $\mu$s 18 $\mu$s
The FLOPs necessary to compute the sparse attention are about 20$\times$ lower than for the full attention. The reduction of runtime is 65%, because the computation is not as well optimized as FlashAttention and includes overheads. Note that end-to-end timings do not only depend on the budget, but also on how balanced the clusters are and whether the heads all perform the same number of FLOPs. This is due to the parallelization over GPUs and over computation cores that we use (see Section 4).
The training takes around 25 minutes per head on one GPU. This is not a limitation since the training is meant to be re-used across contexts and prompts.
6.4 Comparison with the state of the art
We compare Saap with RetrievalAttention (Liu et al., 2024a). For this comparison, we employ a Llama 3 model fine-tuned by Gradient AI to support 262k contexts. We set $\ell$ = $32$, yielding a selectivity of around 5% for textual content.
Table 3 shows that the baseline numbers are often different, probably because of slightly different experimental settings (e.g. prompts). However, the gap between Saap and the full attention baseline is very similar to the RetrievalAttention results. Note that in our experimental setting, the dense part of the attention is 1+2047 tokens, while it is 128+512 for RetrievalAttention. However, for long-range tasks, this does not matter much as it is still a tiny fraction of the attention size (see the ablations in Section 6.5). The critical advantage of Saap is that the slowest stage (training) is performed offline, while RetrievalAttention’s RoarGraph index (Chen et al., 2024a) has to be built once the context is available, which takes several minutes for a 100k context. As mentioned in Section 2, this makes graph-based indexing of a KV cache attractive only in settings where the same KV cache is reused multiple times. In contrast, for Saap, pre-filling the index takes a few seconds on modern GPUs.
6.5 Ablation experiments
We have evaluated numerous more complex variants that did not yield consistent improvements: increasing the model capacity, training a key classification model, fine-tuning the query classification model to the current prompt, using causal masking in the training batches, training on book summarization data rather than books to force long-range queries. In general, we observed that good results from experiments at the level of one head do not guarantee an improved end-to-end performance.
Q-model architecture and training.   In Appendix E.1, we analyze some variants of the Q-model training. This analysis shows that using long-range queries-key pairs for training is useful. Using a residual architecture rather than a simple MLP does not help.
Dense context size.   Saap’s default dense context is 2048 tokens (the attention sink of size 1 plus the 2047 most recent keys). We experiment with some other settings for attention sink+dense context: 1+511, 1+63 and 128+512. The results in Appendix E.2 show that the performance remains similar.
Batched queries.   Saap combines all four queries belonging to the same group into a single search, see Section 5. These queries could also be performed independently. In Appendix E.3, we show that batching the queries corresponding to the same KV head outperforms ranking the buckets independently for each query.
Number of clusters $C$ and probes $\ell$.   We discuss the trade-off in Appendix E.4. The selectivity is roughly proportional to $\ell/C$. Large $C$ provides better quality but is limited for efficiency reasons.
6.6 Limitations
Our method involves an additional training stage. While its computational cost is negligible compared to the LLM training, it involves some expertise (1) to ensure that the model quality is not significantly impacted for the target use-case(s), (2) to reset the hyper-parameters if the ones we report are not adapted for the target model. Another limitation is that our method is a solution for inference, not to training and not for prefill (although we do not see any major obstacle to adopt it for prefill).
7 Conclusion
We introduced Saap, a new method for non-exhaustive attention computations in LLMs. Saap is data adaptive, while incurring negligible run-time or memory overhead compared to exhaustive attention. Our paper has focused on the generation stage, where memory I/O has the strongest impact on performance. We have shown the efficiency of Saap for typical long-context tasks, with significant computation speedups and a small impact on accuracy.
Appendices
In the following appendices, we present a few additional results that complement the main paper. Appendix A provides details regarding the training procedure. Appendix B examines a few LSH-based partitioning methods. Appendices C and D report experimental results for RULER and additional ones for Infinitybench. Appendix E reports the results of ablation experiments. Appendix F details how the Saap algorithm is implemented on GPU.
Appendix A Training details
The training was performed on 40 prompts of 500k tokens each. For these we have access to all key and query vectors from all layers and heads.
Key assignment.   we train the k-means for 10 iterations with a random initialization. This is standard setting for partition-based indexing methods, we use the IVF implementation of the Faiss library (Douze et al., 2024).
Q-model.   We train the Q-model for 40k iterations. At each iteration, we sample 1000 queries vectors from one random prompt. The query vectors are sampled among queries for which the nearest key in the prompt is more than 1024 tokens away in the sequence, to favor long-range queries (around 21% the queries have this property). All the key vectors from the same prompt are used for training, except the attention sink.
The loss of Equation 1 is optimized with the Adam optimizer and a learning rate of $10^{-5}$. Note that we set a high number of iterations to get the best possible setting and because the model’s training time is not critical for the pre-fill or inference speed.
Appendix B Partitioning the KV-cache with LSH
Figure 5: MSE vs. selectivity for three LSH variants, compared to a random selection and a partition based on k-means (the baseline for Saap), on 4 representative heads.
Locality Sensitive Hashing refers to multiple algorithms inspired by the Johnson-Lindenstrauss lemma. In the context of approximate nearest neighbor research, the term is overloaded, as it covers several algorithms and variants. One common component is that the vector to hash $x\in\mathbb{R}^{d}$ is projected onto $r$ directions drawn randomly, which yields the $r$-bit bucket number:
$B(x)=\sum_{i=0}^{r}2^{i}\mathbf{1}[x^{\top}\pi_{i}>0]\in\{0,\dots,2^{r}-1\}.$ (2)
Multiple hash tables.
MagicPig (Chen et al., 2024b) employs several (up to $L=200$) hash tables. The vector search relies on collisions between the buckets: a key is considered only if it is hashed together with the query in at least two hash tables. Using such a large number of hash tables is classical for LSH, since a single hash table is very unbalanced. The theoretical guarantees of LSH apply only when the number of hash tables is large.
Therefore, it is not strictly speaking a partitioning method of the keys. Besides, the large number of hash tables means that the memory used for that indexing structure is larger than the KV-cache itself ($8004L=$ 800 bytes per vector to store, compared with the $2\times 2\times 128=512$ bytes to store the key-value pairs). This is why MagicPig stores the hash tables in CPU RAM.
A single hash table?
In the KDEformer method of Zandieh et al. (2023), a single hash table is considered, with a large $r$, so that the buckets are sparse. The bucket numbers corresponding to they keys $B(k)$ are ordered by their Gray codes [^1] https://en.wikipedia.org/wiki/Gray_code (which the authors re-invent). Then the sequence of keys ordered in this way is partitioned in equal sized bins. At query time, only the bin the query vector “falls” in is visited.
Since successive binary strings ordered by Gray codes differ by only 1 bit, the hope is that nearby buckets end up being nearby in this ordering. However, it is not guaranteed that if $B(k)$ and $B(k^{\prime})$ differ by only 1 bit, then they will be nearby in this enumeration. Indeed, the bit that they differ by is not necessarily the one that changes in successive Gray codes. Besides, it is also unlikely that many keys differ by only one bit because if $r$ is large then most vectors will differ by more than one.
Alternative hash functions.
In the Reformer of Kitaev et al. (2020), the hashing function is replaced with the maximum dot product of random directions:
$B(x)=\underset{i=1..2^{r}}{\mathrm{argmax}}\leavevmode\nobreak\ \leavevmode\nobreak\ x^{\top}\pi_{i}.$ (3)
Compared to Equation (2), each random projection only generates a single bucket instead of a bit for the bucket number. This is close to our k-means approach, if the centroids were drawn randomly.
Similar to KDEFormer, the sequence of keys is sorted and split arbitrarily, but the bucket numbering does not have a significance here. The authors observe that it is beneficial to use multiple hash tables.
Discussion.
The theoretical grounding of these LSH based methods is fragile. They rely on a supposed property explicited in the HyperAttention work (Han et al., 2023):
A useful property of this LSH variant is that its buckets are ordered in such a way that geometrically adjacent buckets have consecutive buckets.
There is an error in this reasoning: it is not possible to map a high-dimensional space to a sequence while maintaining neighborhood relations; otherwise, nearest neighbor search in high dimensions would be as easy as in 1D.
Besides, the subsequent arbitrary splitting of linear sequences into regular buckets defined to minimize computations defeats the purpose of the original data buckets.
Experiments.
We compare different LSH approaches in the tradeoff between selectivity and mean squared error (MSE) for a few attention heads. The MSE is computed between the approximate attention output and the exact one. We use random sampling as the baseline. Since the attention mechanism can be seen as Gaussian kernel smoothing Zhang et al. (2024a), random sampling is actually a reasonable estimator, that reaches MSE=0 when all vectors are sampled (we use sampling without replacement).
All the variants have a parameter that adjusts the tradeoff: for KDEFormer and Reformer, fewer larger buckets improve the accuracy and increase the computational cost (1 bucket covering all keys and queries is equivalent to full attention); for MagicPig’s LSH with multiple hash tables, increasing the number $L$ of hash tables improves the coverage of the key vectors.
Figure 5 shows that for this metric, the KDEFormer and Reformer are on-par or slightly worse than random sampling. Note, however, that these two paritioning methods were developed to be used at pre-fill time, so their performance could be better in a self-attention setting where the contexts embeddings are matched among themselves. In fact, Reformer uses the query corresponding to a key to assign that key to a bucket instead of attempting to assign the key itself. MagicPig implements the classical multi-hashtable LSH, so it performs better, but note that since MagicPig uses multiple hash tables, it does not yield a (single) partition of the dataset. The k-means approach obtains much better results, mainly because it is data-adaptive.
These LSH based methods are confronted with the same bucketing balancing issues that we encounter in Saap. MagicPig resolves this balancing issue by running the search on CPU, that is more tolerant to irregular computation patterns than GPUs. The two other variants arbitrarily split the keys and queries into fixed-size buckets that a handled easily by block computations. Saap attempts to balance the buckets at training time with a specific loss term.
Appendix C RULER results
In Table 4, we report the results of the RULER (Hsieh et al., 2024) subtasks: multi-key, multi-query and VT for two context lengths 65k and 128k. Surprisingly, the SAAP scores are often above the full attention scores. This can be explained by the fact that the sparse attention focuses on only a relevant subset of the prompts, thus it has a denoising effect. SAAP scores are also higher than k-means in three out of five tasks.
Table 4: Scores with a Llama 3.1-8B model using rope extrapolation to increase the context length of 500k and different partitioning schemes. We report in bold the best results among approximate attention schemes. Each subtask has the corresponding context size in parentheses.
RULER (Hsieh et al., 2024)
Methods multi-key multi-query VT
context length 65k 128k 65k 128k 65k
Full attention 39.0 25.8 54.5 62.7 50.5
StreamingLLM (1+2047) 3.6 2.0 1.9 1.7 5.0
K-means (roped inputs, $\ell=32$) 27.6 14.2 78.4 60.7 78.4
K-means (no rope, $\ell=32$) 34.4 24.2 65.0 67.7 54.2
Saap, K-means+Q ($\ell=32$) 46.6 31.8 70.0 65.7 85.0
Appendix D InfiniteBench results
Additional results.
Figure 6 complements  Figure 4 and illustrates the task-dependent behaviors described in Section 6.2. . The results are more noisy, yet Saap does perform well for most tasks.
Figure 6: Performance on selected InfiniteBench tasks
Appendix E Ablation experiments
In these experiments we evaluate our method with variations of a few key parameters.
E.1 Variations on architecture and training setting
Figure 7 shows the performance on the NIH and En.QA tasks using different architectures. It is comparable to Figure 4 with a few more variants for the Q model:
  • “Q-model short-range” is the standard 2-layer MLP, but trained without filtering out short-range queries ;
  • “Q-model residual” relies on a residual architecture with the same capacity as the MLP, also trained without filtering out short-range queries.
The NIH plot shows that training on long-range queries only improves the accuracy significantly. Besides, the residual architecture is not an improvement over the MLP. The En.QA results are more contrasted but the task is more noisy, so we chose to standardize on the MLP architecture with short-range filtering.
Figure 7: Comparison of architectural variants on the needle-in-haystack (left) and InfiniteBench En.QA (right) tasks.
E.2 Dense context size
Figure 8 shows the performance of various dense context sizes. Depending on the task, some dense context settings are more efficient. We chose 1+2047 that works well on the needle-in-haystack task.
Figure 8: Comparison of dense context sizes on the needle-in-haystack (left) and InfiniteBench En.QA (right) tasks. The dense context is reported as “128+512” means that the first 128 tokens of the sequence and the most recent 512 ones are accessed exactly (dense).
E.3 Batched queries
Figure 9 shows the effect of batching the 4 queries belonging to the same KV head.
  • "Batched": Jointly rank partitions across multiple queries to identify the top-k partitions that we want to access.
  • "Independent": Access the top-k ranked partitions for each query separately.
Interestingly, we observe that the batched setting often outperforms the independent setting, which ought to be more data-adaptive.
Figure 9: Comparison of batched and independent queries on the needle-in-haystack (left) and InfiniteBench En.QA (right) tasks.
E.4 Guidance for the number of clusters and the number of visited clusters
We carried out additional experiments with a varying number $C$ of partitions in a setting similar to Figure 5: we measure the MSE (approximation error) with k-means partitioning (for a few arbitrary layers and heads). We vary the number of visited clusters $\ell$ for several settings of $C$. The selectivity roughly is proportional to $\ell/C$.
Figure 10 shows that with a fixed $\ell/C$ and selectivity, a lower (better) MSE can be obtained with a larger number of clusters $C$. This is also a classical observation in partition-based indexing (Douze et al., 2024). However, a larger $C$ poses several problems: (1) the assignment operation is more expensive, (2) it is harder to train – especially for a Q-model, (3) the GPU implementation becomes less efficient because the memory regions to access are more fragmented. Therefore, we set $C$=1024 by default.
Selectivity is a monotonously increasing function of the hyperparameter $\ell$, the number of clusters visited per query. Table 5 and Figure 11 show that depending on the task, the selectivity increases faster with $\ell$ , eg. for NIH the selectivity at $\ell=32$ is around 4% while it is 48.5% for Math.F. This high selectivity can be explained by the fact that the Math.F benchmark is built on long sequences of numbers, which is a different domain compared to natural language. For this task Saap is ineffective, because at this selectivity level, the speedup with respect to full attention is not significant.
By contrast, for Code.D the selectivity is close to that of natural language task en.QA. This is interesting because, while the LLM training data includes code, the key and query partitioning models did not: in this case Saap is relatively robust.
The relationship between $\ell$ and selectivity can also be observed in the curves of Figure 4, where the score vs. selectivity operating points are obtained by varying $\ell=4,8,...,256$.
Figure 10: Error (MSE) as a function of the number of clusters $C$ and the number of probes $\ell$.
Table 5: Selectivity when varying the number $\ell$ of visited clusters.
$\ell$
balanced
NIH 128k
NIH 500k
Retr.N
Retr.P
Retr.KV
Math.F
Code.D
En.QA
En.MC
4 0.4 0.5 0.8 1.8 1.5 3.7 3.2 0.7 0.4 0.4
8 0.8 1.0 1.5 3.7 3.3 7.0 7.4 1.5 0.9 0.9
16 1.6 2.0 3.0 8.1 7.2 14.1 21.4 3.0 2.0 2.0
32 3.1 4.0 5.3 17.4 15.9 28.7 48.5 5.4 4.1 4.1
64 6.2 7.8 9.2 36.7 35.3 60.3 73.5 9.5 8.2 8.2
128 12.5 15.1 16.2 82.6 83.0 87.7 88.4 16.5 16.2 16.2
Figure 11: Selectivity at various choices for $\ell$, the number of clusters visited.
Appendix F Algorithm details
In this section, we will describe the algorithm details of the Saap Triton kernel. Given a query $Q\in\mathbb{R}^{1\times d}$, the 1+2047 keys and values that are accessed exactly are $K_{l},V_{l}\in\mathbb{R}^{N_{l}\times d}$; the remaining keys and values are $K_{s},V_{s}\in\mathbb{R}^{N_{s}\times d}$, Every $K_{s},V_{s}$ has a corresponding partition assignment $A_{i}\in\{1,...C\}$ for $i\in\{1,...,N_{s}\}$, where $C$ is the number of buckets.
@@ Appendix F Algorithm details
The key and value vectors are stored in HBM (high-bandwidth memory) that is slower than the on-chip SRAM.
For a given query $Q$, we assign the $\ell$ most promising buckets in $B\in\{1,...,C\}^{\ell}$
We parallelize the program by launching $\ell+1$ CUDA blocks, where the first block will handle the 1+2047 $K_{l},V_{l}$, and each of the remaining blocks will handle a given bucket and will perform the sparse attention. Our kernel assumes the grouped-query attention setting where keys and values share the same head, with only the queries having $n_{h}$ independent heads. Given this setting, we avoid duplicating the $K,V$ heads and we compute the $n_{h}$ heads from $Q$ together, by leveraging the fact that the sequence length for $Q$ is 1, so that the computation is equivalent to transposing the head and the sequence dimension for $Q$.
Algorithm 1 Partial attention kernel PAttn
0:  Inputs $Q,O^{*}\in\mathbb{R}^{n_{h}\times d}$, $K,V\in\mathbb{R}^{B_{c}\times d}$, $m^{*}\in\mathbb{R}^{n_{h}\times 1}$ and $s^{*}\in\mathbb{R}^{n_{h}\times 1}$ all in on-chip SRAM
  Compute $s:=QK/\sqrt{d}$
  Compute $m_{i}:=\text{rowmax}(m^{*},A)$
  Update $O^{*}:=O^{*}\cdot\exp(m^{*}-m_{i})+V\exp(s-m_{i})$
  Update $s^{*}:=s^{*}\cdot\exp(m^{*}-m_{i})+\exp(s-m_{i})$
  Update $m^{*}:=m_{i}$
Algorithm 2 Inverted list computation kernel
0:  Inputs $Q\in\mathbb{R}^{n_{h}\times d}$, $K_{s},V_{s}\in\mathbb{R}^{N_{s}\times d}$, $K_{l},V_{l}\in\mathbb{R}^{N_{l}\times d}$, buckets to visit: $B\in\{1,...,C\}^{\ell}$, Off and Idx, key block size $B_{c}$, and CUDA block id $b_{i}$.
  Load $Q$ from HBM to on-chip SRAM
  Initialize accumulator $O_{b_{i}}:=0_{d}$
  Initialize sumexp accumulator $s_{b_{i}}:=0$
  Initialize max accumulator $m_{b_{i}}:=-\infty$
  if $\ell>=b_{i}>0$ then
     Read $c=B[b_{i}-1]$, the bucket id to visit
     Let $I^{c}$ be the indices of Idx that correspond to bucket $c_{i}$ by indexing $I^{c}=\texttt{Idx}[\texttt{Off}[c]:\texttt{Off}[c+1]]$
     Divide the indices $I^{c}$ in $T=\lceil\frac{|I^{c}|}{B_{c}}\rceil$ blocks $\{I^{c}_{1},...I^{c}_{T}\}$ of size $B_{c}$
     for $0\leq j<T$ do
        Load indices $I_{j}^{c}$ to SRAM.
        Strided load $K_{s}^{j},V_{s}^{j}:=K_{s}[I_{j}^{c}],V_{s}[I_{j}^{c}]$ to SRAM
        $\texttt{PAttn}(Q,K_{s}^{j},V_{s}^{j},m^{b_{i}},s_{b_{i}},O_{b_{i}})$
     end for
  else if $b_{i}==0$ then
     compute block size $T=\lceil\frac{N_{l}}{B_{c}}\rceil$
     for $0\leq j<T$ do
        compute indices $i_{0}=B_{c}j$ and $i_{1}=B_{c}(j+1)$
        $\texttt{PAttn}(Q,K_{l}[i_{0}:i_{1}],V_{l}[i_{0}:i_{1}],m^{b_{i}},s_{b_{i}},O_{b_{i}})$
     end for
  end if
  Write $O_{b_{i}},s_{b_{i}},m_{b_{i}}$ to HBM
  In a separate kernel, for every $n_{h}$ compute merge-attention:
  Load $O_{0},s_{0},m_{0}$
  Set $m^{*}:=m_{0}$, $O=O_{0}$ and $s=s_{0}$
  for $1\leq i\leq\ell$ do
     Load $O_{i},s_{i},m_{i}$
     Compute $\hat{m}:=\max(m^{*},m_{i})$
     if $m_{b_{i}}<m^{*}$ then
        Compute $\alpha:=\exp(m_{i}-\hat{m})$
        Update $O_{i}:=\alpha O_{i}$
        Update $s_{i}:=\alpha s_{i}$
     else
        Compute $\alpha:=\exp(m^{*}-\hat{m})$
        Update $O:=\alpha O$
        Update $s:=\alpha s$
     end if
     Set $m^{*}:=\hat{m}$, $s:=s+s_{i}$ and $O:=O+O_{i}$
  end for
  return  $O/s$
Once the partial results are computed, we apply the merge-attention function from xFormers (Lefaudeux et al., 2022) to aggregate the partial results into the final output.